find all the angles between 0degree to 360degree which satisfy the equation:
sin2x=-0.7880

Answer by nerdybill(6950)

find all the angles between 0degree to 360degree which satisfy the equation:
sin2x=-0.7880
.
Let y = 2x
then
instead of
sin2x=-0.7880
you would have
sin y =-0.7880
.
Take the inverse sin (arcsin) of both sides giving you
y = -52 deg
but
y = 2x
-52 = 2x
-26 deg = x (quadrant four)
or
360-26 = x
334 deg = x
.
since we are looking for all the angles between 0degree to 360degree:
it also falls on the third quadrant:
180+52 = 2x
232 = 2x
116 deg = x
.
Answer: 116 and 334 degrees